Click Here to receive Service Notes automatically via email.
Date | Description |
---|---|
03/01/2006 | TruView will be restarted at approximately 6:00 a.m. on 3/1/2006. The restart is necessary to apply some configuration changes needed following the recent Banner upgrade. |
02/28/2006 | For today only the PureMessage Quarantine Digests will be processed at 5:00 pm instead of 3:00 pm. |
02/27/2006 | The email server was queueing rather than delivering incoming and outgoing mail. This has been corrected and the queue is clearing now. All mail should be delivered soon. |
02/27/2006 | All network services have been restored for Ophelia Hall. Preliminary analysis shows a strand of backbone fiber optic cable to Ophelia Parrish has failed. We have switched to a backup fiber optic cable. |
02/24/2006 | TruView will be restarted during the scheduled maintenance window on Friday. The maintenance window begins at 8:00 p.m. This restart is needed for some configuration changes related to the recent Banner upgrade.
The restart has been cancelled pending more testing. |
02/22/2006 | A few faculty and staff had problems with network printing. The cause of the problem was identified as a corrupt file generated from the Banner system. This file is critical to several internal IT processes. Banner support has been contacted. We have restored network printing for the affected individuals. We are continuing to review all other systems and processes that may have been affected by the bad data. |
02/16/2006 | The email server has experienced a slow down in delivering of mail (incoming and outgoing). Mail not delivered immediately has been queueing on the server. The server slow down has been corrected and mail the mail queue should be cleared soon. |
02/16/2006 | A network switch failed in Centennial Hall. The switch is under maintenance and has been called in to the vendor. A replacement switch is scheduled to arrive by 5 PM Thursday. 48 network ports in Centennial Hall were affected by this problem. We've restored temporary network service for 44 of the 48 network ports. |
02/15/2006 | Internet email services have been restored. |
02/10/2006 | At 7:31 pm on Thursday 2/9/2006, the Equitracs core accounting service failed. During this time printing was not impacted. To restore the service a manual restart of the service was necessary. This was completed at approximately 10:15 am. We apologize for any inconvenience this may have caused. We are examining the possible cause of this failure. |
02/08/2006 | All login attempts to TruView between 10:54 a.m. and 11:09 a.m. on 2/8/06 failed due to an error on a dependent webserver. The cause of the failure is being investigated. Some departmental websites may not have been available at this time as well. All services have been restored. |
02/08/2006 | There was a network outage on the Email Ticket system around 5:00 PM Feb 7th. The network had been restored by 9:00 AM this morning to the system. The system handles the processing of emails to helpdesk@truman.edu into Help Desk tickets. All emails that were sent after 5:00 PM Feb 7th have now been processed. |
02/06/2006 | The electrical power outage in Ophelia Parrish has been corrected. One of the network switches was severely damaged with both power supplies destroyed. We've made temporary repairs to restore network service. All network services are back up in Ophelia Parrish. |
02/03/2006 | Truman is presently experiencing an Internet disruption. MOREnet our Internet Service Provider is experiencing disruption of service from Kansas City to Columbia. They are investigating the problem and hope to have it resolved soon.
This disruption does not impact our internal network, thus on-campus communications are operational. Network drives, internal email, and local resources are available. If you have further questions, please contact the IT Services Help Desk at 785-4544 or via e-mail at helpdesk@truman.edu |
02/03/2006 | An internet outage has been reported. Technical Services it examining this report. The local network is functioning, we are investigating our connection to MOREnet. Our outside internet service provider. |
02/03/2006 | Email delivery for email entering and leaving Truman has been resumed. We are continuing to research the cause of the problem. |
02/01/2006 | The computer network at the University Farm is currently down due to a problem with CableOne. CableOne did a cable system upgrade last night that is causing cable service outages for many residents in the University Farm area. CableOne is working on the problem and they will notify us when cable services are restored. |
01/31/2006 | FYI. The old Windows Student domain was shutdown and removed from service this morning. |
01/30/2006 | The WebMail email interface to the old faculty and staff email system is back up. The problem was caused by a problem between the Gold web server and PHP. The problem has been corrected. |
01/26/2006 | Off-campus SSH access to the Gold server has been restricted due to a SSH attack. If you are off-campus and requrie SSH access to Gold please use the Truman VPN client. That will allow you to SSH to Gold from off-campus. |
02/06/2006 | We're continuing to work with the vendor on several printer accounting related issues. We will resume full printer accounting services as soon as possible. |
01/24/2006 | The Student print server SDC has failed. We are making an emergency switch to a new student print server named TP1 to restore student printing services. TP1 hosted student printing services are now back up in classrooms and labs for Windows based systems. We are continuing to work on restoring Mac printing services. ITS web documentation for residence hall users will be updated shortly. |
01/23/2006 | Network services were restored for Centennial Hall fourth floor north. This occurred at approximately 9:00 p.m. A power cycle of all equipment on the fourth floor was required. This will be looked into further by the Network Services team. |
01/23/2006 | Disruption to network services for Centennial Hall 4th floor north was reported at approximately 7:30 p.m.. Network services is responding to the call. |
01/15/2006 | Krypton was not processing print jobs and could not be connected to via RDC. System had to be rebooted manually. This seems to happen often on weekends. The last several reboots are followed by first, a pop-up comment about trouble starting "Web Trends" and then a pop-up message that "at least one service failed to start." |